Evaluate the integral:

∫(t to 0) (3si+(3s^2)j+17k)ds
1
Expert's answer
2020-09-21T14:52:26-0400
"\displaystyle\int_{0}^t(3s\vec{i}+3s^2\vec{j}+17\vec{k})ds="

"=\displaystyle\int_{0}^t3sds\ \vec{i}+\displaystyle\int_{0}^t3s^2ds\ \vec{j}+\displaystyle\int_{0}^t17ds\ \vec{k}="

"=[\dfrac{3s^2}{2}]\begin{matrix}
 t\\
 0
\end{matrix}\ \vec{i}+[\dfrac{3s^3}{3}]\begin{matrix}
 t\\
 0
\end{matrix}\ \vec{j}+[17s]\begin{matrix}
 t\\
 0
\end{matrix}\ \vec{k}="

"=\dfrac{3t^2}{2}\ \vec{i}+t^3\ \vec{j}+17t\ \vec{k}"

"\displaystyle\int_{0}^t(3s\vec{i}+3s^2\vec{j}+17\vec{k})ds=\dfrac{3t^2}{2}\ \vec{i}+t^3\ \vec{j}+17t\ \vec{k}"


"\displaystyle\int_{t}^0(3s\vec{i}+3s^2\vec{j}+17\vec{k})ds=-\dfrac{3t^2}{2}\ \vec{i}-t^3\ \vec{j}-17t\ \vec{k}"